40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3239     Entry time: Fri Jul 16 16:12:31 2010
Author: Alberto 
Type: Configuration 
Category: Computers 
Subject: c1susvme1/2 rebooted 

Today I noticed that the FE SYNC counters of c1susvme1/2 on the RFM network screen were stuck at 16384. I tried to reboot the machines to fix the problem but it didn't work.

The BS watchdog tripped off when I did that, because I had forgotten to disable it. I had to wait for a few minutes before it settled down again.

Later I also re-locked the mode cleaner. But before I could do it, Rana had to reduce the MC_L offset for me.

ELOG V3.1.3-